# Designing Guiding Principles for NLP for Healthcare: A Case Study of Maternal Health

### 19 Dec 2023 v1

@@
ABSTRACT
Objective: An ethical framework for the use of large language models (LLMs) is urgently needed to shape how natural language processing (NLP) tools are used for healthcare applications. Drawing directly from the voices of those most affected, we propose a set of guiding principles for the use of NLP in healthcare, with examples based on applications in maternal health.
Materials and Methods: We led an interactive session centered on an LLM-based chatbot demonstration during a full-day workshop with 39 participants, and additionally surveyed 30 healthcare workers and 30 birthing people about their values, needs, and perceptions of AI and LLMs. We conducted quantitative and qualitative analyses of the interactive discussions to consolidate our findings into a set of guiding principles.
Results: Using the case study of maternal health, we propose nine principles for ethical use of LLMs, grouped into three categories: (i) contextual significance, (ii) measurements, and (iii) who/what is valued. We describe rationales underlying these principles and provide practical advice.
Discussion: Healthcare faces existing challenges including the balance of power in clinician-patient relationships, systemic health disparities, historical injustices, and economic constraints. Our principles serve as a framework for surfacing key considerations when deploying LLMs in medicine, as well as providing a methodological pattern for other researchers to follow.
Conclusion: This set of principles can serve as a resource to practitioners working on maternal health and other healthcare fields to emphasize the importance of technical nuance, historical context, and inclusive design when developing LLMs for use in clinical settings.
1 OBJECTIVE
As generative natural language processing (NLP) methods and large language models (LLMs) have increased in size and performance, so has excitement increased about their clinical use [34]. However, due to the sensitivity of health-related data and the potential for harmful negative outcomes, employing NLP technologies in clinical settings requires careful consideration. Healthcare involves numerous stakeholders who can provide useful contributions towards the design of these technologies, e.g., care-seekers, clinicians, researchers, and others all have a shared interest in the development and use of these technologies.
@@ 1 OBJECTIVE
In this work, we identify a set of guiding principles that support well-grounded and ethical uses of NLP technologies in healthcare, developed with input from diverse stakeholder groups. We employ a participatory design framework [27] to amplify the voices of stakeholder groups who have previously been underrepresented in the discussion around use of NLP technologies in healthcare.
@@
We develop our guidelines by focusing on a case study of maternal healthcare in the United States. We choose maternal health for three reasons. First, pregnancy and childbirth are common events that often comprise a person’s sole or major interaction with the healthcare system. Second, there are many prior research studies and applications of NLP methods focused on maternal healthcare [23, 17, 21, 35]. Third, maternal health is a “perfect storm” of healthcare vulnerabilities, with historical biases and power dynamics influencing care [12, 25]. Studying maternal health allows us to map our investigation and recommendations over a set of specific and grounded opportunities and risks.
@@
We solicited key considerations about an LLM-based chatbot and NLP tools from diverse stakeholders, including clinicians, birthing people, researchers, community health workers, government and non-profit workers, and others. We collected perspectives in two main ways (Figure 1). First, we introduced stakeholders to NLP technologies, allowing them to surface concerns, opportunities, and discussion through interaction with these technologies in a controlled setting. Second, we conducted surveys to gather information about questions, priorities, and risks.
@@
Through analysis of the collected data, iterative rounds of discussion, and literature review, we consolidated and identified nine guiding principles for the use of NLP in healthcare, organized according to three themes: context, measurements, and values. These themes are inspired by our maternal health case study but provide a transferable blueprint for mapping tensions in other healthcare areas, supporting NLP practitioners and healthcare administrators on decisions related to the design and deployment of NLP tools.
2 BACKGROUND AND SIGNIFICANCE
Ethical guidelines for NLP/AI and healthcare
@@ 2 BACKGROUND AND SIGNIFICANCE
Prior work has developed guidelines and frameworks for the use of artificial intelligence (AI) for healthcare, and provide important guidance for AI practitioners. These works take a general view of healthcare and are often focused on ML practitioners and their processes. For example, Chen et al. [9] and Wiens et al. [37] provide an overview of the NLP development pipeline and recommendations focused on each pipeline step. The recommendations in these works are based on literature reviews and healthcare examples.
@@
We build on these works by focusing our investigation on a specific healthcare context and direct feedback from its stakeholders, then drawing broader lessons for the rest of healthcare. In particular, we elicit values from various stakeholders using a framework introduced by Jakesch et al. [20], which aims to model the “human process” at the center of healthcare [16]. As recommended by Rajkomar et al. [30] and following prior work [36], we follow research process that draws on participatory design, a collaborative design approach that involves stakeholders and end users in the design process for new technologies [27], such that their needs drive design decisions and result in tools that better address their concerns and challenges [14, 22, 31, 19]. We also make use of participatory design in the formation of our guiding principles; uniquely, we survey stakeholders connected to a specific healthcare topic to better illuminate tensions and priorities for NLP usage for healthcare.
Maternal health: vulnerabilities and NLP applications
@@
Rates of pregnancy-related deaths and complications have increased over the last thirty years in the U.S., with particularly high rates among Hispanic and Black birthing people and non-U.S. citizens [5, 11, 12, 24, 25]. For example, Black women are three times more likely to die in childbirth than white women [12, 25]. Most frustratingly, 46% of maternal deaths of Black women and 33% of maternal deaths of white women are estimated as preventable [6], but measurements of these problems are challenging [13].
@@
Such urgent problems need every tool in the toolbox, and NLP models can facilitate the analysis of large-scale datasets. Prior work has used NLP methods to study the prevalence of adverse outcomes via social media data [23], extract lactation information from drug labels [17], predict high-risk pregnancies from electronic health records (EHR) [21], examine how news media discuss pregnancy and exposure to isotretinoin [26], and assist in analysis of phone interviews about breastfeeding [35]. In our review of this prior work in NLP, we observe the following patterns: (i) many studies attempt to surveil and predict psychological states of the birthing person, e.g., predicting postpartum depression, and (ii) most studies use either social media or EHR data, with a bias in studies published at NLP venues towards social media data produced by birthing people.
3 MATERIALS AND METHODS
@@
In Figure 1, we provide an overview of our methods, which center around a structured interface for participants to query an LLM-based chatbot. After participants have interacted with this demonstration, they move to virtual breakout discussion sessions led by volunteer moderators, and a survey, in which we asked participants to express their perceptions of NLP tools for maternal healthcare. Non-workshop participants wrote individual responses to our discussion questions rather than worked together with a moderator. Our study was approved by the IRB at the Allen Institute for AI.
@@
Figure 1: Study overview, including the three participant cohorts, the chatbot demonstration, and the surveys and discussions followed by analysis and design of the guiding principles.
3.1 Participant Cohorts
We ran this study with three participant cohorts. Using the Prolific [^1] https://www.prolific.com/ survey platform, we recruited one cohort of birthing people (gave birth in the last five years) and one cohort of healthcare workers (have ever worked in any healthcare profession). We required that all participants be over 18 years old and be located in the U.S.
For the third cohort, we led a live, virtual session with a group of participants who received training about LLMs and their risks. Our session was part of a virtual workshop hosted by the Association of American Medical Colleges (AAMC) on NLP and maternal health in which participants were introduced to the basics of NLP, applications of NLP to maternal health, and biases in NLP tools. Most but not all of these participants also worked in healthcare, though compared to our cohort from Prolific, these participants more often worked in community and research roles rather than as clinicians.
@@ 3.1 Participant Cohorts
For full descriptions of these cohorts, see Table 1 in §4.1.
3.2 Design of chatbot demonstration
@@
We built an interface wrapper (Figure 2) around GPT-3.5 from OpenAI. [^2] Compared to GPT-4, which was also accessible at the time of our study, GPT-3.5 is less expensive and has lower rate limits, which was especially important for the workshop setting, where 100+ people were interacting with the bot at any moment. Our goal was to demonstrate high performance for the bot rather than create a reproducible system; otherwise, an open model would have been preferable. Our goal with this demonstration was to give participants a guided but unconstrained interaction with a chatbot, gathering their queries, structured feedback, and perceptions related to the chatbot and maternal health. For many of our participants (see Figure 6 in the Appendix), this was their first direct interaction with a modern LLM. We included prominent warnings about (i) hallucinations and unreliability, (ii) data collection (see Figure 2), and (iii) the importance of always seeking the advice of a healthcare professional. We also provided examples of the kinds of maternal health questions one might ask a chatbot.
Figure 2: Screenshots of our chatbot demo. Responses were generated using GPT-3.5 from OpenAI. Users could ask multiple questions about maternal health. The workshop demonstration additionally included a free-text entry box for participants to add descriptions of situations in which they might use the chatbot.
3.3 Design of discussion sessions
Workshop participants
@@ 3.3 Design of discussion sessions
Workshop participants were divided into eight breakout rooms and participated in a 30-minute free-form discussion session, moderated by workshop organizers and trained volunteers. Discussion sessions were recorded after obtaining consent from all participants. During the session, participants were shown the following discussion prompts:
@@
We received video and audio recordings for four out of eight sessions as well as written comments from five sessions, which we included in our analysis.
Healthcare workers and birthing people
@@
Non-workshop participants first interacted with our chatbot demonstration, then wrote answers to our discussion questions as part of a followup survey (like the workshop participants).
3.4 Survey design
@@
Our goals in the survey were to (i) elicit participants’ perceptions of NLP tools, (ii) learn about participants’ information seeking goals, and (iii) build collaborative rankings of values that should guide NLP uses for maternal health. Surveys given to each cohort were identical except that for birthing people, questions about career were replaced with questions about whether the participant would like their healthcare team to use and/or disclose their use of AI tools. Figure 3 provides a condensed view of the survey we gave to participants; see Appendix C for the full set of survey questions. Prolific workers were paid an average of \$15/hour.
Figure 3: Simplified view of the survey given to participants. The question about trust is drawn from Baughan et al. [3], while the questions about general AI attitudes are drawn from a frequently reused set by Nickell and Pinto [28]. The listed values were taken from Jakesch et al. [20]; definitions were provided for each value. Questions in the final box were shown only to the cohort of birthing people.
4 RESULTS

### 23 Jan 2024 v2

# NLP for Maternal Healthcare: Perspectives and Guiding Principles in the Age of LLMs

@@
Abstract.
This is an unpublished pre-print and has not been peer-reviewed. Before citing, please check authors’ websites for updated version and publication information. Comments welcome.
Ethical frameworks for the use of natural language processing (NLP) are urgently needed to shape how large language models (LLMs) and similar tools are used for healthcare applications. Healthcare faces existing challenges including the balance of power in clinician-patient relationships, systemic health disparities, historical injustices, and economic constraints. Drawing directly from the voices of those most affected, and focusing on a case study of a specific healthcare setting, we propose a set of guiding principles for the use of NLP in maternal healthcare. We led an interactive session centered on an LLM-based chatbot demonstration during a full-day workshop with 39 participants, and additionally surveyed 30 healthcare workers and 30 birthing people about their values, needs, and perceptions of NLP tools in the context of maternal health. We conducted quantitative and qualitative analyses of the survey results and interactive discussions to consolidate our findings into a set of guiding principles. We propose nine principles for ethical use of NLP for maternal healthcare, grouped into three themes: (i) recognizing contextual significance (ii) holistic measurements, and (iii) who/what is valued. For each principle, we describe its underlying rationale and provide practical advice. This set of principles can provide a methodological pattern for other researchers and serve as a resource to practitioners working on maternal health and other healthcare fields to emphasize the importance of technical nuance, historical context, and inclusive design when developing NLP technologies for clinical use.
maternal health, natural language processing, large language models, ethical guidelines
copyright: noneccs: Applied computing Health informaticsccs: Computing methodologies Natural language processingccs: Social and professional topics Medical technologies
As natural language processing (NLP) methods and large language models (LLMs) have increased in size and performance, so has hype and excitement increased about their clinical use (Thirunavukarasu et al., 2023). Recent work experimenting with generative LLMs has found promising results in comparison to physicians, including for tasks such as addressing public health concerns (Ayers et al., 2023b), answering care-seekers’ questions (Ayers et al., 2023a), and engaging in diagnostic dialogues (Tu et al., 2024). But healthcare and ethics researchers have also highlighted the safety risks (Mehandru et al., 2022), biases (Xiao et al., 2023), inaccuracies (Boag et al., 2021), and other problems arising from NLP tools applied to healthcare contexts, as well as general risks of LLMs (Bender et al., 2021; Weidinger et al., 2022). Given the sensitivity of medical data and the potential for harmful negative outcomes, deciding when and how to employ NLP technologies in clinical settings requires careful consideration.
Two key challenges complicate these decisions. First, numerous stakeholders have a shared interest in the development and use of these technologies; care-seekers, clinicians, researchers, hospital administration, and other groups can all provide design and decision-making input, but these voices are not equally represented, may conflict, and are rarely brought together during decisions around system design and implementation. Second, most research developing ethical guidelines has studied high level applications of machine learning for healthcare, across (i) a range of medical topics and (ii) across a range of machine learning methods. There is a gap for research grounded in a specific healthcare setting where risks, benefits, and stakeholder perceptions can be fully explored, focusing specifically on NLP technologies and the new risks posed by LLMs.
@@ Abstract.
In this work, we focus on a case study of maternal healthcare in the United States. We employ a participatory design framework (Muller and Kuhn, 1993) to amplify the voices of diverse stakeholder groups who have previously been underrepresented in the discussion around use of NLP technologies in healthcare. Drawing on the input from these groups, we identify a set of guiding principles that support well-grounded and ethical uses of NLP and LLM technologies in maternal healthcare.
@@
We choose maternal health for three reasons. First, there are many prior research studies and applications of NLP methods focused on maternal healthcare (Lee et al., 2021; Goodrum et al., 2019; Knowles et al., 2014; Valdeolivar-Hernandez et al., 2023). Second, pregnancy and childbirth are common events that often comprise a person’s sole or major interaction with the healthcare system, increasing the significance and also abundance of perspectives on this topic. Third, maternal health is a “perfect storm” of healthcare vulnerabilities, with historical biases and power dynamics influencing care; for example, maternal health in the U.S. has received much attention in recent years due to the high morbidity and mortality of birthing people and significant racial inequities (Creanga et al., 2015; Martin and Montagne, 2017). Focusing on maternal health allows us to map our investigation of NLP applications over a set of specific and grounded opportunities and risks.
@@
We solicited key considerations about an LLM-based chatbot and NLP tools from diverse stakeholders, including clinicians, birthing people, researchers, community health workers, government and non-profit workers, and others. We collected perspectives in two main ways (Figure 1). First, we introduced stakeholders to NLP technologies, allowing them to surface concerns, opportunities, and discussion through interaction with these technologies in a controlled setting. Second, we conducted post-interaction surveys to gather information about priorities, opportunities, and risks.
@@
Through analysis of the collected data, iterative rounds of discussion, and literature review, we consolidated and identified nine guiding principles for the use of NLP in maternal healthcare, organized according to three themes: context, measurements, and values. These themes are inspired by our maternal health case study but provide a transferable blueprint for mapping tensions in other healthcare areas, supporting NLP practitioners and healthcare administrators on decisions related to the design and deployment of NLP tools.
1. Related Work
1.1. NLP and LLMs for Healthcare
NLP methods have been applied to many kinds of health data, both in research and in industry applications. Traditionally, NLP for healthcare has focused on structured tasks like information extraction from clinical notes (Savova et al., 2010; Wang et al., 2018; Si et al., 2019) and outcome prediction (Shamout et al., 2020; Naik et al., 2022; van Aken et al., 2021), or on public health surveillance of social media (Nikfarjam et al., 2015; Glaz et al., 2019; Baclic et al., 2020). Technical advances in LLMs have shifted this focus, with recent work applying generative NLP tools to a wide variety of question-answering use cases, including engaging in diagnostic dialogues (Tu et al., 2024), addressing medical scenarios (Nori et al., 2023; Ueda et al., 2023), and generating responses for patients asking questions on social media sites (Ayers et al., 2023a). Research has also focused on developing large datasets (Fleming et al., 2023) or studying clinician adoption of LLMs (Sivaraman et al., 2023; Kim et al., 2023). On the applied side, as one example, Epic Systems, the leading provider of electronic health record (EHR) systems in the United States, is bringing GPT-4 into EHRs to help clinicians communicate with patients (Turner, 2023). And in practice, everyday users are already exploring their healthcare questions with generic LLM-based chat tools; according to one analysis of a publicly-released dataset of user-GPT conversation histories, approximately 3% of user queries are health-related (Ouyang et al., 2023).
These use cases have fueled both optimism (Vaishya et al., 2023; Lee et al., 2023) and scrutiny (Wornow et al., 2023; Vaishya et al., 2023). Within the fairness, accountability, and transparency (FAccT) community, researchers have studied many general risks and challenges arising with LLMs (Bender et al., 2021; Weidinger et al., 2022) and also highlighted risks specific to healthcare settings. For example, recent work on ethical uses of machine learning for healthcare has studied bias measurement (Xiao et al., 2023), explainability methods (Panigutti et al., 2020; Balagopalan et al., 2022; Poulain et al., 2023), dataset documentation gaps (Rostamzadeh et al., 2022), and procedures for ethical implementations and deployments (Sendak et al., 2020). Compared to research on other application areas, these works emphasize the particular privacy issues associated with clinical data (Xiao et al., 2023), the importance of optimizing for the right outcome variables (health outcomes rather than institutional costs) (Obermeyer and Mullainathan, 2019), and interactions between healthcare workers and AI systems (Fogliato et al., 2022) that require explainability (Panigutti et al., 2020; Balagopalan et al., 2022; Poulain et al., 2023).
Most of these works addressing healthcare contexts study machine learning methods more broadly, rather than focusing on NLP methods and the specific challenges arising from new tools like LLMs. We build on these works by focusing on NLP tools for maternal health and by eliciting perceptions from multiple stakeholder groups.
1.2. Ethical Guidelines for NLP/AI and Healthcare
@@
Prior work has developed important guidelines and frameworks for the use of machine learning (ML) and artificial intelligence methods for healthcare. These works take a general view of healthcare and are often focused on ML practitioners and their processes. For example, Mccradden et al. (2023) focuses on ethical guidelines for ML-informed clinical decision-making, and Chen et al. (2021) and Wiens et al. (2019) provide overviews of the ML/NLP development pipeline and recommendations focused on each pipeline step. The recommendations in these works are based on literature reviews and broad sets of healthcare examples. While most of these guidelines consider healthcare broadly, Sendak et al. (2020) design guidelines based on the deployment of a specific sepsis-detection machine learning tool, and Petti et al. (2023) focused on developing ethical guidelines for the use of NLP and AI methods for early detection of Alzheimer’s disease. Our work builds most directly on the latter study, as we also focus on linguistic data and NLP methods.
@@
We build on these works by combining in a novel way the following goals: (i) focusing on NLP and new LLM technologies, (ii) grounding our work in maternal health as a specific healthcare context, and (iii) soliciting direct feedback from multiple stakeholder groups. In particular, we elicit values from various stakeholders using a framework introduced by Jakesch et al. (2022), which aims to model the “human process” at the center of healthcare (Ghassemi and Mohamed, 2022). As recommended by Rajkomar et al. (2018) and following prior work (Vedam et al., 2019), we follow a research process that draws on participatory design, a collaborative design approach that involves stakeholders and end users in the design process for new technologies (Muller and Kuhn, 1993), such that their needs drive design decisions and result in tools that better address their concerns and challenges (Delgado et al., 2023; Laura Ramírez Galleguillos and Coşkun, 2020; Sorensen et al., 2023; Inie et al., 2023). We engage with these stakeholder groups not to “solve” issues with NLP tools but to unearth concerns and themes (Sloane et al., 2022). We also make use of participatory design in the formation of our guiding principles; uniquely, we survey stakeholders connected to a specific healthcare topic to better illuminate tensions and priorities for NLP usage for healthcare.
1.3. Maternal Health: Vulnerabilities and NLP Applications
@@
The U.S. is experiencing an urgent and worsening maternal mortality crisis. Rates of pregnancy-related deaths and complications have increased over the last thirty years in the U.S., with particularly high rates among Hispanic and Black birthing people and non-U.S. citizens (Berg et al., 2010; Creanga et al., 2014, 2015; MacDorman et al., 2016; Martin and Montagne, 2017). For example, Black women are three times more likely to die in childbirth than white women (Creanga et al., 2015; Martin and Montagne, 2017). Most frustratingly, 46% of maternal deaths of Black women and 33% of maternal deaths of white women are estimated as preventable (Berg et al., 2005), but measurements of these problems are challenging (Declercq et al., 2023). Additional dangers face birthing people, such as postpartum depression (Stewart and Vigod, 2016), and urgently need supportive solutions.
@@
Researchers and healthcare practitioners have sometimes turned to NLP methods to try to address various challenges in the maternal health space. For example, prior research has used NLP methods to study the prevalence of adverse outcomes via social media data (Lee et al., 2021), extract lactation information from drug labels (Goodrum et al., 2019), predict high-risk pregnancies from electronic health records (EHR) (Knowles et al., 2014), examine how news media discuss pregnancy and exposure to isotretinoin (Mohammadhassanzadeh et al., 2020), and assist in analysis of phone interviews about breastfeeding (Valdeolivar-Hernandez et al., 2023). In our review of this prior work in NLP, we observe the following patterns: (i) many studies attempt to surveil and predict psychological states of the birthing person, e.g., predicting postpartum depression, and (ii) most studies use either social media or EHR data, with a bias in studies published at NLP venues towards social media data produced by birthing people.
2. Methods
@@
In Figure 1, we provide an overview of our methods. Participants first engage with a structured interface to query an LLM-based chatbot. After participants have interacted with this demonstration, they move to discussions (either virtual breakout discussion sessions led by volunteer moderators or independent written comments), and a survey, in which we asked participants to express their perceptions of NLP tools for maternal healthcare. Our study was approved by the institutional review board (IRB) at the Allen Institute for AI.
@@
Figure 1. Study overview, including the three participant cohorts, the chatbot demonstration, and the surveys and discussions followed by analysis and design of the guiding principles.
2.1. Participant Cohorts
@@
We ran this study with three participant cohorts. For full demographic descriptions of these cohorts, see Table 1 in §3.1.
2.1.1. Healthcare workers and birthing people
Using the Prolific [^1] https://www.prolific.com/ online survey platform, we recruited one cohort of birthing people (gave birth in the last five years) and one cohort of healthcare workers (have ever worked in any healthcare profession). We required that all participants be over 18 years old and be located in the U.S. Prolific workers were paid an average of \$15/hour.
2.1.2. Workshop participants
For the third cohort, we led a live, virtual session with a group of participants who first received training about LLMs and their risks. Our session was held at the end of a virtual workshop hosted by a U.S.-based medical nonprofit. The workshop was themed around NLP and maternal health equity, and in sessions preceding ours, participants were introduced to the basics of NLP, heard research talks about applications of NLP to maternal health, and learned about biases and ethical challenges in NLP tools. The workshop was open to the public, and participants were solicited through the nonprofit’s email listserv, social media, and word of mouth. Most but not all of these participants also worked in healthcare, though compared to our cohort from Prolific, these participants more often worked in community and research roles rather than as clinicians (see Table 1).
2.2. Design of chatbot demonstration
@@
We built an interface wrapper (Figure 10) around GPT-3.5 from OpenAI. [^2] Compared to GPT-4, which was also accessible at the time of our study, GPT-3.5 is less expensive and has lower rate limits, which was especially important for the workshop setting, where 100+ people were interacting with the bot at any moment. Our goal was to demonstrate high performance for the bot rather than create a reproducible system; otherwise, an open model would have been preferable. Our goal with this demonstration was to give participants a guided but unconstrained interaction with a chatbot, gathering their queries, structured feedback, and perceptions related to the chatbot and maternal health. For many of our participants (see Figure 3 in the Appendix), this was their first direct interaction with a modern LLM. We included prominent warnings about (i) privacy and data collection and (ii) the importance of always seeking the advice of a healthcare professional (see Figure 10 in Appendix B). We also provided examples of the kinds of maternal health questions one might ask a chatbot.
2.3. Design of discussion sessions
2.3.1. Workshop participants
@@
Workshop participants were divided into eight virtual breakout rooms and participated in a 30-minute free-form discussion session, moderated by the authors and trained volunteers. Discussion sessions were recorded after obtaining consent from all participants. During the session, participants were shown the following discussion prompts:
  1. (1)
@@
We collected video and audio recordings for four out of eight sessions as well as written comments (collected via Mentimeter [^3] https://www.mentimeter.com/) from five sessions, which we included in our analysis.
2.3.2. Healthcare workers and birthing people
@@
Like the workshop participants, non-workshop participants first interacted with our chatbot demonstration and then wrote answers to our discussion questions as part of a followup survey. Unlike the workshop participants, non-workshop participants did not participate in discussions with groups or with a moderator but instead completed the study independently.
2.4. Survey design
@@
Our goals in the survey were to (i) elicit participants’ general perceptions of NLP tools, (ii) learn about participants’ information seeking goals, and (iii) build collaborative rankings of values that should guide NLP uses for maternal health. Surveys given to each cohort were identical except that for birthing people, questions about healthcare careers were replaced with questions about whether the participant would like their healthcare team to use and/or disclose their use of AI tools. See Appendix C for the full set of survey questions.
Participants were first asked about their generalized trust then asked about their trust in healthcare providers; the format of these questions is drawn from Baughan et al. (2023). Participants were then asked about their familiarity with NLP tools like ChatGPT, as well as their general perceptions of AI; these questions were taken from a frequently reused set by (Nickell and Pinto, 1986). Next, participants were asked to select five out of ten ethical values that should guide the use of NLP for maternal health. The listed values were taken from Jakesch et al. (2022) in a study of variations in attitudes towards AI by different demographic groups. We provided definitions (also drawn from Jakesch et al. (2022)) and example healthcare applications of NLP systems. Participants were asked about their information seeking needs (where and to whom they turn with maternal health questions) and asked their opinion about the effects of NLP tools on maternal health team members and on their own job (if applicable). Finally, participants answered a series of demographic questions, including specifying any professional role they had ever taken in maternal health (e.g., worked as a midwife).
3. Results
3.1. Survey Results
We release our survey results publicly to support future research. [^4] We remove demographic features for the workshop participants, as these participants interacted with one another and could potentially identify others’ survey responses. Participants consented to the sharing of their responses. All survey data is available in supplementary materials and at a public Github repository. Figure 2 shows how frequently each value was selected by each cohort when asked to “select any five values … that you think are most important for NLP systems for maternal health.” Overall, safety, privacy, and performance were selected more often by birthing people and healthcare workers in comparison to workshop participants, who were more likely to select inclusiveness. The birthing people and workshop participants were united in being more likely to select human autonomy than the healthcare workers, and less likely than the healthcare workers to select transparency and accountability. Overall, we do not find large differences in the value selections within cohorts across stratifications for generalized trust, trust in healthcare workers, or generalized AI perceptions (see Appendix A).
Figure 2. Frequencies of selected values. Each participant was asked to select five values from a list of 12 curated by Jakesch et al. (2022). Participants were given definitions of the values, also drawn from Jakesch et al. (2022). Birthing people and healthcare workers overall responded more similarly to each other than to the workshop participants, but this was not always the case (e.g., transparency, human autonomy).
Table 1 provides a summary of participant demographics and background. The workshop participants tended to represent the non-profit, community health, research, and governmental sectors in contrast to clinicians such as physicians and nurses that were more common in the Prolific cohort of healthcare workers. The workshop participants represented a more even distribution across racial/ethnic groups, but all cohorts had little representation in the East Asian, South Asian, and Southeast Asian groups. [^5] We attempted to correct this balance on Prolific by supplementing our results with an additional cohort of East Asian, South Asian, and Southeast Asian participants, but we found that the Prolific survey platform included very few such participants who also met our other criteria (e.g., gave birth in the last five years). Overall, workshop participants reported less experience with AI and NLP than birthing people and healthcare workers recruited from Prolific (see Appendix A).